OpenStudy (anonymous):

A hexagon is inscribed in a circle. If the difference between the area of the circle and the area of the hexagon is 36 m2, use the formula for the area of a sector to approximate the radius r of the circle. (Round your answer to three decimal places.)

OpenStudy (anonymous):

A buddy is in the same class and we both had trig in high school, so we are wondering how everyone else is doing since we're having so much trouble. I managed to finally figure out how to solve using sectors only it requires knowing the relationship between the sides of a 30-60-90 triangle, which is a couple chapters down the road. This is my work: \[A=\frac{1}{2}r^{2}θ\] \[A=12rh\] \[\frac{1}{2}r^{2}θ−\frac{1}{2}rh=36\] \[\theta = \frac{\pi}{3}\] \[\frac{1}{2}r^{2}(\frac{\Pi}{3})−\frac{1}{2}rh=36\] \[h=\frac{\sqrt{3}}{2}r\] \[\frac{1}{2}r^{2}(\frac {\Pi}{3})−(\frac{1}{2})r(\frac{\sqrt{3}}{2})r=36\] \[\frac{1}{2}r^{2}(\frac{\Pi}{3})−\frac{1}{2}(\frac{\sqrt{3}}{2})r^{2}=36\] \[\frac{1}{2}r^{2}[(\frac{\Pi}{3})−(\frac{\sqrt{3}}{2})]=36\] \[6(\frac{1}{2}r^{2}[(\frac{\Pi}{3})−(\frac{\sqrt{3}}{2})])=36\] \[3r^{2}[(\frac{\Pi}{3})−(\frac{\sqrt{3}}{2})])=36\] \[r^{2}[(\frac{\Pi}{3})−(\frac{\sqrt{3}}{2})])=12\] \[r=\sqrt{12 \div [(\frac{\Pi}{3})−(\frac{\sqrt{3}}{2})])}\] \[r \approx 8.139\] While this works, is there a way to do it with trigonometric functions, but still apply sector theory?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (akashdeepdeb):

It is a regular hexagon right?

OpenStudy (anonymous):

Yeah, that's assumed from the problem given, so each of the six triangles formed from the hexagon would be equilateral triangles with 60 degree angle measures.

OpenStudy (akashdeepdeb):

Yes excatly! :D

OpenStudy (akashdeepdeb):

And then the area of the circle would be pi*r*r - area of hexagon pi*r*r - [3 sqrt 3*r^2/2] = 36 Just find this and you'll get radius!! :D

OpenStudy (anonymous):

That's what I did with my work shown above. I was wondering if there was a way to do it without involving any roots and by using trig functions because our book technically hasn't covered 30-60-90 triangles yet.

OpenStudy (akashdeepdeb):

Yes you can do it but this method is the EASIEST and the fastest. And where would you require 30-60-90 triangles? :D

OpenStudy (anonymous):

I did it when I split the equilateral triangle in half to find h in terms of r and got h = (sqrt3/2)r

OpenStudy (akashdeepdeb):

Hahah...But why? :') You already have a direct formula for equilateral triangle = sqrt 3*a^2/4 Why do you want to slpit it? XD

OpenStudy (anonymous):

. . . I like right triangles :D I usually end up breaking things down too far and over-analyze them. It was making a lot of sense to break it that way even if its the long way.

OpenStudy (akashdeepdeb):

Hahahaha!! LOL But you understood this now right? :D

OpenStudy (anonymous):

Not really, I remember getting that equation before and I was wondering how the pi*r*^2 part worked because the formula for the area of a sector has a measure of theta involved.

OpenStudy (dan815):